# Gradient Masking All-at-Once: Ensemble Everything Everywhere Is Not Robust

### v1

@@ Abstract
In this short note, we show that this defense is not robust to adversarial attack. We first show that the defense’s randomness and ensembling method cause severe gradient masking. We then use standard adaptive attack techniques to reduce the robust accuracy of the defense from $48\%$ to $1\%$ on CIFAR-100 and from $62\%$ to $4\%$ on CIFAR-10, under the $\ell_{\infty}$ norm threat model with $\varepsilon=8/255$.
@@ 1 Introduction
Unfortunately, we show that this defense is not robust and that the existing evaluation does suffer from significant gradient masking. We build a stronger adaptive attack that significantly reduces the robust accuracy to just $1\%$ for CIFAR-100, or $4\%$ for CIFAR-10.111A similar attack was concurrently shown in [23].
In summary, while the approach proposed in [9] remains an interesting and novel way to yield informative visualizations and image manipulations, it does not yield a defense that is robust to worst case perturbations.
@@ 4 Breaking the Defense
We use the official codebase released by the authors [^3] https://github.com/stanislavfort/ensemble-everything-everywhere to train the defense on CIFAR-10 and CIFAR-100. Our models achieve clear accuracy of $88.9\%$ on CIFAR-10, and $64.1\%$ on CIFAR-100. We then run the AutoAttack evaluation considered by the authors, and obtain robust accuracies of $61.8\%$ on CIFAR-10 and $47.9\%$ on CIFAR-100 (see Table 1). This is consistent with the results reported by the authors in their repository. [^4] The paper [9] reports slightly higher robustness numbers using larger models, longer training runs, and combining with adversarial training. As these models and their training code are not publicly released, we rely on the official training code. Since this defense relies heavily on randomness, we conduct 10 runs for all results and then report the mean and standard deviation.
Our attack.
Our attack is rather simple: the main trick is to attack the defense without the CrossMax operator, which causes gradient masking. This, together with some standard tricks, suffices to reduce robust accuracy to below random chance. We summarize our attack steps below:
@@
  • Transfer from a model without CrossMax: we found that if we replace the CrossMax aggregation by a simple mean, we can reliably reduce the model’s robust accuracy. This suggests that either (1) all the robustness comes from the CrossMax, or (2) the CrossMax causes gradient masking. We find that the latter is true: if we take the adversarial examples optimized for the model with a mean aggregation, and then simply transfer these to the target model, the robust accuracy is reduced by about 1/3131/31 / 3.555A similar approach was concurrently described in [23].
@@
[…]
@@
+ finetune $19.8\pm 2.0$ $12.5\pm 1.9$
+ restarts $\mathbf{3.5\pm 0.8}$ $\mathbf{0.3\pm 0.5}$
@@
Figure 2: For different attack strategies, we plot the attack’s loss averaged over 100 examples: (1) directly optimizing the target model (in green) has trouble converging due to gradient masking; (2) attacking the “source” model with a mean aggregation works well; (3) transferring the attack from the source model to the target model (red) outperforms the attack that directly optimizes over the target model; (4) once we are close to an adversarial example, we switch to optimizing over the target model (olive) to make the attack succeed.
@@ 5 Discussion
Unfortunately, this defense did not end up giving us a fourth category of approaches for adversarial robustness. We believe there are a few lessons to be learned from this defense and the (cautious) enthusiasm that it originally caused.
@@
  • We could analyze and break this defense using only existing techniques [32, 25, 2]. As already noted a few years ago in [25], the time seems to have passed where breaking a defense requires new research ideas. Now, it is mainly an exercise in rigorously applying known techniques until the defense breaks (or one fails after exhausting all attempts). The issue with current defense evaluations is thus primarily methodological: defense authors should try as hard as possible to cover all possible attack strategies before claiming empirical robustness.
  • Although our attack did not need any new techniques, it does highlight (again) the importance of performing careful adaptive evaluations. Although it is tempting to run standard off-the-shelf evaluation tools such as AutoAttack [8], this is not sufficient. As we show above, AutoAttack maxes out at only a $52\%$ attack success rate on CIFAR-100, while our adaptive attack achieves $>99\%$ success.
  • Identifying these types of flaws requires the use of specialized tools designed specifically for this purpose. The original paper attempts to argue robustness, in part, by showing that the defended models have interpretable gradients. This is because many robust models indeed have interpretable gradients [27]. But the converse is not necessarily true [10]—as we showed again in the case of this model. Its gradients are highly perceptually aligned, but the model exhibits almost no robustness to adversarial attacks.

### v2

# Evaluating the Robustness of the “Ensemble Everything Everywhere” Defense

@@ Abstract
In this short note, we show that this defense’ robustness to adversarial attack is overclaimed. We first show that the defense’s randomness and ensembling method cause severe gradient masking. We then use standard adaptive attack techniques to reduce the robust accuracy of the defense from $48\%$ to $14\%$ on CIFAR-100 and from $62\%$ to $11\%$ on CIFAR-10, under the $\ell_{\infty}$ norm threat model with $\varepsilon=8/255$.
An earlier version of this report by Zhang, Nikolić, Carlini and Tramèr contained incorrect attack results due to a bug that created perturbations 10×10$\times$10 × larger than the bound allowed by the defense [10]. This new version corrects this bug and the resulting claims about the defense’s robustness. We also add results from a concurrent robustness evaluation by Schlarmann, Croce and Hein [26].
@@ 1 Introduction
Unfortunately, we show that this defense is not as robust as initially claimed and that the existing evaluation does suffer from gradient masking. We build a stronger adaptive attack that reduces the robust accuracy to $14\%$ for CIFAR-100, or $11\%$ for CIFAR-10.
In summary, while the approach proposed in [11] is an interesting and novel way to yield informative visualizations and image manipulations, it does not yield a defense that achieves high robustness to worst case perturbations.
These results are reminiscent of a previous defense—ME-Net [35]—which also yields interpretable perturbations but was shown to have only low robustness on CIFAR-10 under adaptive attacks [28].
@@ 4 Evaluating the Defense’s Robustness
We use the official codebase released by the authors [^2] https://github.com/stanislavfort/ensemble-everything-everywhere to train the defense on CIFAR-10 and CIFAR-100. Our models achieve clean accuracy of $88.9\%$ on CIFAR-10, and $64.1\%$ on CIFAR-100. We then run the AutoAttack evaluation considered by the authors, and obtain robust accuracies of $61.8\%$ on CIFAR-10 and $47.9\%$ on CIFAR-100 (see Table 1). This is consistent with the results reported by the authors in their repository. [^3] The paper [11] reports slightly higher robustness numbers using larger models, longer training runs, and combining with adversarial training. As these models and their training code are not publicly released, we rely on the official training code. Since this defense relies heavily on randomness, we conduct 10 runs for all results and then report the mean and standard deviation. In contrast to [11], we attack all samples, even if they are classified incorrectly already in clean evaluation. Since the forward pass of the target model is subject to randomness, such samples are not necessarily always classified incorrectly. Thus it is important to attack also these samples [16].
Our PGD attack.
Our attack is rather simple: the main trick is to attack the defense without the CrossMax operator, which causes gradient masking. We summarize our attack steps below:
@@
  • Transfer from a model without CrossMax: we found that if we replace the CrossMax aggregation by a simple mean, we can reliably reduce the model’s robust accuracy. This suggests that either (1) all the robustness comes from the CrossMax, or (2) the CrossMax causes gradient masking. We find that the latter is true: if we take the adversarial examples optimized for the model with a mean aggregation, and then simply transfer these to the target model, the robust accuracy is reduced by about 1/3131/31 / 3.444This was also described in [26].
@@
  • Bag of Tricks: A standard PGD+EoT attack transferred from a model without CrossMax works reasonably well, and reduces robustness to around 20–30%. We then add further simple tricks—additional PGD steps (up to 400), more EoT steps (up to 100), a change of loss function (the Hinge loss of [6] instead of a standard cross-entropy)—to reduce the robustness further to around $11.3\%$ on CIFAR-10 and $13.8\%$ on CIFAR-100.
@@
+ bag of tricks $11.3\pm 2.5$ $13.8\pm 2.1$
APGD w/ transfer, EoT $16.4\pm 2.2$ $16.8\pm 1.3$
+ larger initial radius $15.3\pm 2.7$ $16.0\pm 1.5$
+ 1000 total steps $13.8\pm 2.5$ $15.9\pm 1.7$
Best-of-both $\mathbf{10.9\pm 2.4}$ $\mathbf{13.6\pm 1.4}$
@@
Figure 2: (a) For different attack strategies, we plot the attack’s loss averaged over 100 examples: (1) directly optimizing the target model (in green) has trouble converging due to gradient masking; (2) attacking the “source” model with a mean aggregation (blue) works well; (3) transferring the attack from the source model to the target model (red) outperforms the attack that directly optimizes over the target model. (b) We show an ablation on the amount of EoT iterations with the APGD attack on 32 CIFAR-10 samples using 100 steps. We observe that it is crucial to use a sufficiently high amount of iterations, however, there is no consistent improvement beyond 100 EoT iterations.
Our APGD attack.
This attack is based on APGD [8], which is known to be a strong attack due to its adaptive step size schedule. We maximize the cross-entropy loss, using 100 steps by default. To adapt this attack to the target model at hand, we use the following approach:
  • Transfer from a model without CrossMax: Similar to our PGD attack, we utilize transfer from a model that replaces CrossMax with mean aggregation.
  • Expectation over Transformation (EoT): We use 100 EoT iterations. An ablation on the amount of EoT iterations is shown in Figure 2.
  • Larger initial radius: We use a larger perturbation budget at the beginning of the attack as proposed in [9]. This allows the attack to find promising adversarial directions more easily. The perturbation radius is then gradually decreased during the attack according to the following schedule: $3\epsilon$ for $30\%$ of total iterations, $2\epsilon$ for $30\%$ of total iterations, and $\epsilon$ for $40\%$ of total iterations. This is an existing functionality of APGD that is by default deactivated.
@@ 5 Discussion
Unfortunately, this defense did not end up giving us a fourth category of approaches for (high) adversarial robustness. We believe there are a few lessons to be learned from this defense and the (cautious) enthusiasm that it originally caused.
@@
  • We could analyze and partially break this defense using only existing techniques [36, 28, 2]. As already noted a few years ago in [28], the time seems to have passed where breaking a defense requires new research ideas. Now, it is mainly an exercise in rigorously applying known techniques until the defense breaks (or one fails after exhausting all attempts). The issue with current defense evaluations is thus primarily methodological: defense authors should try as hard as possible to cover all possible attack strategies before claiming empirical robustness.
  • Although our attack did not need any new techniques, it does highlight (again) the importance of performing careful adaptive evaluations. Although it is tempting to run standard off-the-shelf evaluation tools such as AutoAttack [8], this is not sufficient. As we show above, AutoAttack maxes out at only a $52\%$ attack success rate on CIFAR-100, while our adaptive attack achieves $\geq 85\%$ success.
  • Identifying these types of flaws requires the use of specialized tools designed specifically for this purpose. The original paper attempts to argue robustness, in part, by showing that the defended models have interpretable gradients. This is because many robust models indeed have interpretable gradients [30]. But the converse is not necessarily true [12]—as we showed again in the case of this model. Its gradients are highly perceptually aligned, but the model exhibits only very small robustness to adversarial attacks. A similar phenomenon had been observed with a prior defense—ME-Net [35]—which uses matrix estimation techniques to yield interpretable gradients, but only very moderate adversarial robustness [28].
  • Attacks have to be rigorously tested too, and adaptive evaluations can contain their own bugs! The implementation in the original version of this report stacked multiple attack stages, and incorrectly computed the $\ell_{\infty}$ bound for each stage with respect to the output of the previous stage. As a result, after $k$ stages the adversarial perturbations could be as large as $k\cdot 8/255$. (see [10] for details.
    A good practice to avoid such errors (and which we used in our amended PGD attack) was to implement the attack in a script that saves the final adversarial examples to disk, and then have a second standalone script that solely evaluates the attack against the original defense and data.
    We also used this script to visualize adversarial examples (see Figure 3). While some of these showcase interpretable perturbations (as claimed in the original defense paper), others do not seem to.
(a) CIFAR-10
(b) CIFAR-100
Figure 3: Examples of clean images (top row) with the adversarial examples found by our attack (middle row), and the corresponding perturbation (bottom row, centered and magnified). In some cases, the perturbations are clearly interpretable, e.g, CIFAR-10 “cat” $\to$ “dog” (2nd from right) or CIFAR-100 “sea” $\to$ “apple” (5th from left).
@@
Acknowledgments.
We thank Stanislav Fort for discovering the bug in the original implementation of the attack of Zhang et al., and for valuable discussions about the defense and comments about this write-up.